OpenStudy (anonymous):

how to factor x^2+9x=20+8x

OpenStudy (nubeer):

x^2+9x-8x-20= 0 x^2+x-20 = 0 x^2 +5x-4x-20= 0 x(x+5) -4 (x+5) =0 (x+5)(x-4) =0